OpenStudy (anonymous):

cos(2x+10)=sin(x=20)

OpenStudy (raden):

use the identity : sin(x) = cos(90-x)

OpenStudy (raden):

wait, in right side is that sin(x+20) ?

OpenStudy (anonymous):

yes

OpenStudy (raden):

ok, look at this how to solve this

OpenStudy (raden):

cos(2x+10)=sin(x+20) cos(2x+10)=cos(90 -(x+20)) cos(2x+10)=cos(90 - x - 20) cos(2x+10)=cos(70 - x ) cancel out the cos, in both side 2x+10 = 70 - x now, can u solve for x

OpenStudy (anonymous):

thank you!

OpenStudy (raden):

you're welcome but what's ur answer ?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

??

OpenStudy (raden):

x = ............. ?

OpenStudy (raden):

that's right, good finishing ^_^